OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1) medal! @SithsAndGiggles

OpenStudy (kirbykirby):

P(1) is the expression with 1 term: LHS: 2 RHS: 1(1+1) = 1(2) = 2 P(k) .. just replace all n's with k's: LHS: 2 + 4 + 6 + ... + 2k RHS: k(k+1) P(k+1), here you find the (k+1)-th term, so: LHS: 2+ 4 + 6 + ... + 2k + 2(k+1) = 2+ 4 + 6 + ... + 2k + (2k+2) RHS: (k+1)[(k+1)+1] = (k+1)[k+2]

OpenStudy (anonymous):

thanks! @kirbykirby

OpenStudy (anonymous):

wait what do lhs and rhs mean?

OpenStudy (kirbykirby):

left hand side = LHS right hand side = RHS I just used those terms to designate which side of the equality I was working on

OpenStudy (anonymous):

ok i dont understand

OpenStudy (anonymous):

is that all there is?

OpenStudy (kirbykirby):

If you have the equation 3x = 2, then LHS = 3x RHS = 2

OpenStudy (anonymous):

i know

OpenStudy (kirbykirby):

well yes that's pretty much it!

OpenStudy (anonymous):

okay thanks